Two dice are rolled at random. Obtain the probability that the sum of the numbers appearing on the dice is equal to 12.
Answer #1

Total number of sample space when rolling two dice at random= n(S)= 6*6=36

Let E be the event that the sum of number on dice is 12

So, E={6,6}

So, Number of event that the sum of number on dice is 12 = n(E)= n[{1,2}]=1

So, probability that sum of number appearing on dice is 12= P( E)= n(E)/n(S)= 1/36=0.0278